### v2

# ACEGEN: Reinforcement learning of generative chemical agents for drug discovery

Abstract
In recent years, reinforcement learning (RL) has emerged as a valuable tool in drug design, offering the potential to propose and optimize molecules with desired properties. However, striking a balance between capabilities, flexibility, reliability, and efficiency remains challenging due to the complexity of advanced RL algorithms and the significant reliance on specialized code. In this work, we introduce ACEGEN, a comprehensive and streamlined toolkit tailored for generative drug design, built using TorchRL, a modern RL library that offers thoroughly tested reusable components. We validate ACEGEN by benchmarking against other published generative modeling algorithms and show comparable or improved performance. We also show examples of ACEGEN applied in multiple drug discovery case studies. ACEGEN is accessible at https://github.com/acellera/acegen-open and available for use under the MIT license.
[…]
1 Introduction
Drug design is a complex process that involves the identification of biomolecules that have an optimal balance of multiple properties, such as potency, selectivity, bioavailability, and toxicity. In recent years, a diversity of generative modeling solutions have been proposed as a promising approach to partially automate the process of proposing new molecules that simultaneously improve multiple desired properties in the design-make-test-analysis cycle 1, 2. These models typically employ machine learning algorithms to generate molecular candidates, but it remains challenging to efficiently search the vast chemical space to identify molecules with optimal properties 3, which is so large that it is not practically enumerable in a naive manner.
Reinforcement learning (RL) has emerged as a possible solution to explore this chemical space 4, 5 with an increasing focus on efficiency 6, 7. RL 8 consists of a family of machine learning algorithms that use feedback as a learning signal. Through a trial and error process, RL algorithms can adapt their search strategy to the specific characteristics of compounds associated with desirable properties, and guide the generative models to propose novel molecules with similar characteristics. These methods use a customized reward function that assigns values to molecules, tailored specifically for the application or research question being addressed. To navigate the molecular space and find the most highly rewarded molecules, RL generative agents alternate between two phases: a data collection phase and a learning phase. During the data collection, molecules are generated by the generative model and scored based on the reward function. In the agent learning phase, the collected data is utilized to enhance the agent’s ability to generate interesting and relevant compounds. By leveraging this feedback loop, the generative model can be continually improved to generate compounds with increasingly desirable properties. This is particularly interesting when the available chemical space cannot be enumerated and filtered using the scoring function, either because the scoring functions are very slow or because the search space is very large.
Currently, available implementations for drug discovery using reinforcement learning (RL) rely heavily on custom code 9, 10. This approach tends to favor redundancy, complexity, and limited efficiency, making it difficult to incorporate a diverse set of solutions. However, the RL community has already developed solutions to address these challenges. TorchRL, a comprehensive RL library 11, offers well-tested, independent state-of-the-art RL components. In this work, we adopt TorchRL’s components as building blocks to assemble efficient and reliable drug discovery agents—a practice already successfully applied in diverse domains such as drone control 12 and combinatorial optimization 13. This approach also fosters algorithmic research by enabling the encapsulation of new ideas within new components, which can seamlessly integrate with existing ones. TorchRL operates within the PyTorch ecosystem 14, ensuring not only high-quality standards but also maintenance over time and future development.
To showcase the advantages of ACEGEN, we implement and evaluate well-known language-based algorithms for drug design. Language models can learn complex patterns in text and generate novel sequences, including molecular structures 15, 16. Studies have focused on utilizing the Simplified Molecular Input Line Entry System (SMILES) 17, a string notation system used to represent chemical structures in a compact and standardized manner. Moreover, several prior studies have demonstrated the benefits of combining generative language models and RL 4, 18, 7. Therefore, we re-implement three REINFORCE-based algorithms: REINFORCE 19, REINVENT 4, and AHC 6. Additionally, we adapt general RL algorithms such as A2C 20 and PPO 21 to our problem setting. The REINVENT and AHC methods utilize some level of experience replay; therefore, we also incorporate experience replay into REINFORCE. Similarly, we test PPOD 22, a PPO-based algorithm adapted for experience replay.
To demonstrate the possible use cases and applications of ACEGEN, we have benchmarked these RL algorithms on a sample efficiency benchmark while considering chemistry, conducted an ablation study to better understand the components of a specific RL algorithm, tested a more challenging drug design relevant objective, and lastly presented RL for molecular constrained generative design. ACEGEN code is accessible open source under the MIT license at https://github.com/acellera/acegen-open to facilitate uptake by the community.
2 Methods
2.1 Reinforcement learning setting
Reinforcement learning tasks are formalized as Markov Decision Processes (MDPs) 8. Formally, an MDP is described by a quintuple, denoted as $\langle S,A,R,P,\rho_{0}\rangle$. Here, $S$ represents the set of all possible states in the problem space, $A$ denotes the set of valid actions available to the agent, and $R:S\times A\times S\to\mathbb{R}$ defines the reward function, which assigns numerical values to transitions from one state to another based on the action taken. The function $P:S\times A\to\mathcal{P}(S)$ indicates the transition probability, where $P(s_{t+1}|s,a)$ represents the probability of transitioning to state $s_{t+1}$ from the current state $s$ under action $a$. Lastly, $\rho_{0}$ signifies the initial state distribution.
This approach naturally extends to the problem of sequentially designing molecules. In this context, a parameterized reinforcement learning (RL) policy, denoted as $\pi_{\theta}$, guides molecule design by successively processing the current state representing a partially built molecule and selecting actions $a_{t}$ to make changes to the molecule’s structure at each step until reaching a terminal state. The reward function evaluates the quality of the molecule, assigning a scalar value either at each generative step or at the end of the process when a complete compound is produced. Within this framework, $\pi_{\theta}(a_{t}|s_{t})$ denotes the probability of the policy function, parameterized by $\theta$, taking action $a_{t}$ in state $s_{t}$. The sequence of actions needed to construct a molecule, often referred to as an episode, is represented by $\tau$. $P(\tau|\theta)$ indicates the probability of a trajectory $\tau$ given the policy parameters $\theta$. Furthermore, $R(\tau)$ represents the cumulative sum of rewards over the trajectory $\tau$.
The goal of RL algorithms is to optimize the parameters of the policy $\pi_{\theta}$ to maximize $R(\tau)$. Different methods within the family of policy gradient methods 8, such as REINFORCE 19, A2C (Advantage Actor-Critic) 20, and PPO (Proximal Policy Optimization) 21, are commonly used for this task.
Additionally, various techniques can aid in training. Reward shaping involves modifying the reward function to provide additional feedback to the agent during training, potentially speeding up learning by providing more informative signals. Experience replay involves storing past experiences in a buffer and randomly sampling from it during training, which can improve sample efficiency and stabilize learning. Moreover, incorporating penalty terms, such as a Kullback-Leibler (KL) divergence loss term, can be utilized to encourage the agent to stay close to a reference or prior policy, helping to maintain stability and control the exploration-exploitation trade-off during training. Finally, ranking and selecting only the best K molecules in each given batch can improve efficiency. ACEGEN integrates all the aforementioned algorithms and additional techniques to form a comprehensive suite of state-of-the-art methods for molecular generation.
2.2 Chemical language generative models
One possible approach to sequentially designing molecules is through string representations. Molecular string representations 23 convert molecular graphs into strings and vice versa, allowing molecules to be encoded using specific syntactical rules. This approach formulates the task of molecular generation as a natural language processing (NLP) 24 problem, and the policy models used to sequentially generate molecules represented as strings are called Chemical Language Models (CLMs) 25. In this context, the resulting action space is discrete, with each action denoted as a token. These actions are integrated by concatenating them to form a joint string representation, producing the next state. Each episode begins with a single special start token and can last for a varying number of steps. The episode ends when the agent chooses another special token called the stop token.
ACEGEN currently provides an environment for language model experimentation including SMILES 17, DeepSMILES 26, SELFIES 27, AtomInSmiles 28 and SAFE 29 grammars, complemented by a user-friendly vocabulary class.  Appendix A A in the appendix illustrates how the vocabulary and environment can be easily created and utilized for data generation in ACEGEN.
In practice, CLMs are first trained unsupervised on a bulk of unlabeled data, aimed at learning to generate valid molecules, also referred to as pre-training, which is achieved through the application of the teacher enforcing method. 30. The pre-trained CLM is the starting policy that can be further trained to optimize a specific objective with RL, similar to what happens in large language models. Moreover, this prior policy can be used as an anchor point from which the new RL policy should not deviate excessively.
ACEGEN currently provides pre-trained models for several architectures. More specifically, we provide GRU 31 policies pre-trained on two datasets: ChEMBL 32 and ZINC 16. We also provide weights for an LSTM 33 policy pre-trained on ChEMBL 32 and for a GPT2 34 pre-trained on the REAL Database REAL lead-like compounds 35.
While our repository currently provides methods to create LSTM 33, GRU 31, GPT2 34 , Llama2 36 and Mamba 37 policies, practitioners can similarly include any other architectures of their choice taking advantage of current and future developments in NLP. Users can pre-train their policies with external code and integrate them without any need to modify the internal scripts of the library. Our repository contains a step-by-step tutorial that explains how to do the integration.
Finally, ACEGEN also provides its own pre-training script to train language models, either from ACEGEN or custom, to generate valid molecules through the teacher enforcing method 30. Our code is engineered to adapt to the available computational resources, whether a single GPU or a distributed setup spanning multiple machines and GPUs. This adaptability allows us to efficiently train on datasets of large size.
2.3 Scoring and Evaluation of Molecules
Scoring functions must reflect real-world drug design scenarios and be flexible enough to apply to a range of drug design challenges. Often, overly simplistic objective functions are used for optimization, such as penalized logP 38, or complex solutions tailored to one particular generative model 9. ACEGEN allows the integration of custom scoring functions by providing the flexibility to define them as Python methods that accept strings and return numerical values. Alternatively, ACEGEN is optionally compatible with the MolScore 39 library, which offers a broad range of drug design relevant scoring functions, diversity filters, support for curriculum learning and also contains a benchmarking mode including MolOpt 10, GuacaMol 3, and others.  Appendix B B in the appendix showcases how both scoring options can be accommodated within our codebase. Additionally, the library offers a detailed tutorial, guiding users through the process of implementing and incorporating custom scoring functions seamlessly into the entire workflow to train agents with them.
2.4 RL Agents Training
ACEGEN provides training for RL agents utilizing the following methods: REINFORCE 19, REINVENT 9, AHC 6, A2C 20, and PPO 21, as well as an adapted version of the PPOD algorithm 22. We deviate from the exact implementation of PPOD by omitting the use of an initial expert demonstration, replaying only episodes with high rewards (instead of episodes with high-value predictions), and employing a fixed amount of replay data per batch. All methods are fully configurable, allowing the use of custom scoring functions and models beyond those already provided by the library.
REINVENT and AHC can be considered extensions of the REINFORCE algorithm. Specifically, REINVENT incorporates reward shaping, experience replay, and a penalty for high-likelihood sequences into the basic REINFORCE framework. Additionally, AHC ranks and selects the top K molecules in each batch of data. We find that using the same reward shaping is not compatible with advantage-based methods, and the high-likelihood penalty term is detrimental. Therefore, for A2C, PPO, and PPOD, we introduce a term to the loss function based on the Kullback-Leibler divergence (KL) between the actor policy and a prior policy. This term serves to penalize the policy for deviating too much from a prior, similar in concept to reward shaping in REINVENT and AHC. Incorporating KL constraints is a common practice in research papers aligning language models with custom reward functions, as seen in examples utilizing human feedback 40, 41, 42.
2.5 De-novo, decorative and fragment-linking generation
In drug discovery pipelines, generating molecules from scratch may not always suffice. To meet diverse requirements, ACEGEN offers multiple sampling modes. PromptSMILES 43 is a simple method enabling constrained molecule generation using models pre-trained solely with teacher enforcement on full SMILES, making it possible to generate molecules while adhering to specific chemical sub-structures. Specifically, in addition to de-novo generation, ACEGEN scripts allow for easy configuration of scaffold decoration and fragment-linking molecule generation modes. Constrained sampling modifies only the data collection behavior, operating independently of all other agent components, in line with the TorchRL philosophy. Tutorials on performing constrained generation are provided within the repository.
3 Results
To showcase some of the use cases of ACEGEN, we have benchmarked multiple methods in chemistry benchmarks, conducted an ablation study on REINVENT, tested a challenging drug design objective, and explored constrained RL with PromptSMILES.
3.1 Benchmarking RL performance
To assess different RL algorithms, we have compared performance on the Practical Molecular Optimization (MolOpt) benchmark 10 as implemented in MolScore 39. This benchmark encompasses 23 distinct tasks associated with different objectives to be optimized within a budget of 10,000 molecules. All algorithms use the same GRU policy model architecture as in the original MolOpt benchmark, implemented in ACEGEN and pretrained on a curated subset of ChEMBL 32. We compared REINFORCE 19, REINVENT with the hyperparameters from the original paper 9, REINVENT-MolOpt with optimized hyperparameters for the MolOpt benchmark 10, and AHC with the hyperparameters from the original paper 6. Note that we utilize experience replay in all of the aforementioned algorithms Additionally, we also test A2C 20, PPO 21 which does not utilize any experience replay, and PPOD 22 which is an adaptation of PPO to accommodate experience replay. All algorithms are ACEGEN implementations, and the hyperparameter values used for all of them are provided in our repository. As a sanity check, we compared our implementation of REINVENT with the implementation used in the MolOpt paper. Our simplified re-implementation achieved better results with faster training. These results are shown in Appendix C C in the appendix.
Table 1: Algorithm comparison for the Area Under the Curve (AUC) of the top 10 molecules on MolOpt benchmark objectives. This metric captures the sample efficiency in identifying 10 desirable molecules with respect to the objective. Each algorithm was run 5 times with different seeds, and results were averaged.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Task REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Albuterol_similarity 0.68 $\pm$ 0.03 0.69 $\pm$ 0.02 0.90 $\pm$ 0.01 0.77 $\pm$ 0.02 0.82 $\pm$ 0.04 0.93 $\pm$ 0.02 0.94 $\pm$ 0.00
Amlodipine_MPO 0.55 $\pm$ 0.01 0.56 $\pm$ 0.01 0.65 $\pm$ 0.06 0.56 $\pm$ 0.01 0.55 $\pm$ 0.01 0.58 $\pm$ 0.03 0.68 $\pm$ 0.02
C7H8N2O2 0.83 $\pm$ 0.01 0.82 $\pm$ 0.03 0.90 $\pm$ 0.02 0.76 $\pm$ 0.04 0.84 $\pm$ 0.04 0.89 $\pm$ 0.01 0.89 $\pm$ 0.03
C9H10N2O2PF2Cl 0.70 $\pm$ 0.02 0.70 $\pm$ 0.02 0.76 $\pm$ 0.03 0.68 $\pm$ 0.02 0.69 $\pm$ 0.03 0.66 $\pm$ 0.02 0.79 $\pm$ 0.02
Celecoxxib_rediscovery 0.63 $\pm$ 0.02 0.64 $\pm$ 0.03 0.77 $\pm$ 0.02 0.72 $\pm$ 0.02 0.73 $\pm$ 0.06 0.65 $\pm$ 0.12 0.82 $\pm$ 0.03
DRD2 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.01 0.98 $\pm$ 0.01 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00
Deco_hop 0.63 $\pm$ 0.00 0.63 $\pm$ 0.01 0.67 $\pm$ 0.01 0.64 $\pm$ 0.01 0.62 $\pm$ 0.00 0.62 $\pm$ 0.01 0.66 $\pm$ 0.02
Fexofenadine_MPO 0.71 $\pm$ 0.01 0.71 $\pm$ 0.00 0.80 $\pm$ 0.03 0.72 $\pm$ 0.00 0.71 $\pm$ 0.00 0.73 $\pm$ 0.00 0.78 $\pm$ 0.01
GSK3B 0.84 $\pm$ 0.01 0.84 $\pm$ 0.02 0.92 $\pm$ 0.02 0.82 $\pm$ 0.01 0.85 $\pm$ 0.02 0.90 $\pm$ 0.02 0.92 $\pm$ 0.02
JNK3 0.75 $\pm$ 0.03 0.75 $\pm$ 0.02 0.85 $\pm$ 0.04 0.75 $\pm$ 0.01 0.74 $\pm$ 0.06 0.80 $\pm$ 0.04 0.87 $\pm$ 0.02
Median_molecules_1 0.26 $\pm$ 0.00 0.24 $\pm$ 0.00 0.36 $\pm$ 0.02 0.24 $\pm$ 0.00 0.31 $\pm$ 0.01 0.33 $\pm$ 0.02 0.35 $\pm$ 0.02
Median_molecules_2 0.22 $\pm$ 0.00 0.22 $\pm$ 0.00 0.28 $\pm$ 0.01 0.24 $\pm$ 0.00 0.25 $\pm$ 0.01 0.25 $\pm$ 0.02 0.29 $\pm$ 0.01
Mestranol_similarity 0.60 $\pm$ 0.03 0.55 $\pm$ 0.04 0.85 $\pm$ 0.07 0.66 $\pm$ 0.04 0.69 $\pm$ 0.07 0.75 $\pm$ 0.15 0.89 $\pm$ 0.05
Osimertinib_MPO 0.82 $\pm$ 0.01 0.82 $\pm$ 0.00 0.86 $\pm$ 0.01 0.83 $\pm$ 0.00 0.81 $\pm$ 0.01 0.82 $\pm$ 0.01 0.84 $\pm$ 0.00
Perindopril_MPO 0.48 $\pm$ 0.01 0.47 $\pm$ 0.00 0.54 $\pm$ 0.01 0.47 $\pm$ 0.01 0.48 $\pm$ 0.00 0.50 $\pm$ 0.01 0.53 $\pm$ 0.00
QED 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00 0.94 $\pm$ 0.00
Ranolazine_MPO 0.70 $\pm$ 0.01 0.69 $\pm$ 0.00 0.76 $\pm$ 0.01 0.70 $\pm$ 0.00 0.74 $\pm$ 0.01 0.73 $\pm$ 0.01 0.75 $\pm$ 0.00
Scaffold_hop 0.80 $\pm$ 0.00 0.79 $\pm$ 0.00 0.86 $\pm$ 0.02 0.80 $\pm$ 0.01 0.80 $\pm$ 0.00 0.80 $\pm$ 0.02 0.84 $\pm$ 0.03
Sitagliptin_MPO 0.34 $\pm$ 0.02 0.33 $\pm$ 0.01 0.38 $\pm$ 0.03 0.33 $\pm$ 0.02 0.39 $\pm$ 0.02 0.32 $\pm$ 0.02 0.39 $\pm$ 0.02
Thiothixene_rediscovery 0.41 $\pm$ 0.01 0.41 $\pm$ 0.00 0.56 $\pm$ 0.04 0.45 $\pm$ 0.02 0.48 $\pm$ 0.04 0.48 $\pm$ 0.06 0.58 $\pm$ 0.09
Troglitazone_rediscovery 0.31 $\pm$ 0.02 0.31 $\pm$ 0.02 0.47 $\pm$ 0.05 0.34 $\pm$ 0.01 0.35 $\pm$ 0.02 0.46 $\pm$ 0.07 0.52 $\pm$ 0.06
Valsartan_smarts 0.03 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 0.03 $\pm$ 0.00 0.03 $\pm$ 0.00
Zaleplon_MPO 0.47 $\pm$ 0.01 0.47 $\pm$ 0.01 0.52 $\pm$ 0.01 0.48 $\pm$ 0.01 0.47 $\pm$ 0.01 0.50 $\pm$ 0.02 0.52 $\pm$ 0.01
Total 13.67 13.60 15.65 13.91 14.27 14.65 15.80
As proposed by the MolOpt benchmark authors, Table Table 1 shows algorithm performance on the AUC of the top 10 molecules as an indication of the sample efficiency of identifying 10 desirable molecules with respect to the objective, a representative figure that might be carried forward to later stages of drug design. This shows that PPOD is state-of-the-art with respect to sample efficiency, followed by REINVENT-MolOpt and then PPO. To achieve a higher level understanding of performance, Table 2 shows a selection of metrics aimed to measure maximum performance, efficiency, and exploration (a common trade-off with exploitation and hence efficiency in RL). These results show that REINVENT-MolOpt achieves maximum performance as measured by the average top 10 molecules, PPOD maximum efficiency as measured by AUC of the top 10 molecules, and AHC maximum exploration as measured by the number of unique compounds generated.
Table 2: Algorithm comparison on a selection of metrics analyzing the performance on the MolOpt benchmark. Valid is a sanity check that measures the proportion of valid molecules generated. Top-10 AUC is a measure of sample efficiency as shown in 1. Top-10 Avg is a measure of the absolute best 10 molecules achieved with the budget. Unique is a proxy for exploration which measures the proportion of unique molecules generated. Basic and target chemistry filters (B&T-CF) are the proportion of molecules after filtering out highly idiosyncratic molecules in general or with respect to the pretraining dataset. The B&T-CF Top-10 AUC and B&T-CF Top-10 Avg are recalculated on this subset while also enforcing that the 10 molecules identified offer diverse solutions (Div). Lastly, B&T-CF Diversity is a proxy measure of exploration via sphere exclusion diversity 44 on this subset. Each algorithm was run 5 times with different seeds, and results were averaged and then summed over the 23 tasks. A perfect score for all metrics is 23.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Valid 21.77 21.78 21.74 21.45 18.93 21.78 21.52
Top-10 Avg 15.85 15.69 17.43 16.09 15.67 15.63 17.10
Top-10 AUC 13.67 13.60 15.65 13.91 14.27 14.65 15.80
Unique 22.28 22.63 13.68 22.68 18.38 9.47 10.21
B&T-CF 14.34 14.74 7.00 13.77 7.82 5.81 5.71
B&T-CF Top-10 Avg (Div) 14.83 14.79 16.06 15.25 14.54 14.60 15.78
B&T-CF Top-10 AUC (Div) 12.85 12.81 14.61 13.11 13.32 13.67 14.60
B&T-CF Diversity (SEDiv@1k) 17.39 18.19 10.10 17.55 14.24 8.47 7.96
3.2 Benchmarking RL performance for practical drug discovery
RL optimizes the policy for the cumulative future rewards as provided by either one scoring function or a sum of multiple scoring functions. However, in practical drug discovery, it is not straightforward to write the precise scoring functions that are needed and how to weigh them into a single scalar. For example, scoring functions can sometimes not produce the desired chemistry, or show clear exploitation loopholes 45, 46. In this setting, a sufficient element of regularization to a prior policy is useful, as the prior policy has learned a chemical space distribution based on a specified training dataset of choice which acts as a representation of desirable chemistry by examples. It is for this reason that methods like REINVENT have been designed heuristically to contain terms that try to enforce regularization instead of entirely using the reward signal.
To account for this, we make a modification to the average and AUC of the top 10 molecules to ensure they are diverse by an ECFP4 Tanimoto similarity of less than 0.35 to each other. Moreover, we use sphere exclusion diversity (SEDiv@1k) of de novo molecules 44 which measures the proportion of molecules needed to describe chemical space in a random sample of 1,000 molecules, as a proxy for exploration. This is more representative than the number of unique molecules that could all reside in a close area of chemical space. Lastly, we remove de novo molecules that do not pass a series of filters. This includes a basic chemistry filter (B-CF): logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in 47. Moreover on the assumption that the training dataset describes desirable chemistry by examples, a target chemistry filter (T-CF): logP and molecular weight within $\mu\pm 4\sigma$ of the training dataset distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). The combination of the basic and target chemistry filter is denoted (B&T-CF).
Re-evaluating RL performance with these new metrics (Table 2) it can be seen that a smaller proportion of molecules pass the chemistry filters for REINVENT-MolOpt, A2C, PPO, and PPOD, as expected due to decreased regularization terms to optimize efficiency. Despite this, REINVENT-MolOpt achieves the highest score for the average and AUC of the top 10 diverse molecules followed by PPOD. As a proxy for exploration, REINVENT contains the highest sphere exclusion diversity followed by AHC. Interestingly, REINFORCE maintains a high proportion of molecules passing the chemistry filters despite no explicit regularization term, outperforming REINVENT in absolute performance and efficiency. Table 3 shows the alternative approach of explicitly including these requirements in the reward signal by the addition of the chemistry filters and a diversity filter to each objective. As expected, including these requirements explicitly in the reward function increases the proportion of molecules passing these filters. However, every algorithm drops in performance as measured by the average and AUC top 10 diverse molecules that pass the filters. In this case showing that better overall score optimization and efficiency (including those that measure the quality of chemistry) is achieved by implicit regularization to the prior policy. This could be a result of the specific implementation of the chemistry filters in the reward resulting in a score of 0 if they do not pass; however, optimal implementation of reward signal can in itself be considered an art. Therefore, simpler reward signals are usually better, and, achieving desirable chemistry by regularization enables the design of a simpler reward signal. The fine-tuning of the optimization algorithm, reward signals and regularization might be target-dependent and goes beyond the scope of this work, but it is enabled by ACEGEN.
MolOpt MolOpt-CF
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC PPOD REINFORCE REINVENT REINVENT-MolOpt AHC PPOD
Valid 21.77 21.79 21.74 21.45 21.52 22.00 21.95 21.65 21.71 21.05
Top-10 Avg 15.85 15.67 17.43 16.09 17.10 14.76 14.54 16.17 14.86 15.18
Top-10 AUC 13.67 13.55 15.65 13.91 15.80 12.94 12.78 14.57 13.00 14.01
Unique 22.28 22.67 13.68 22.68 10.21 22.44 22.83 16.63 22.80 14.13
B&T-CF 14.34 14.70 7.00 13.77 5.71 18.44 18.59 13.62 17.72 11.55
B&T-CF Top-10 Avg (Div) 14.95 14.79 16.06 15.25 15.78 14.51 14.33 15.92 14.68 14.86
B&T-CF Top-10 AUC (Div) 12.91 12.81 14.61 13.11 14.60 12.79 12.66 14.35 12.89 13.74
B&T-CF Diversity (SEDiv@1k) 17.39 18.23 10.10 17.55 7.96 16.23 17.65 7.97 17.33 7.92
Table 3: Comparison of algorithm performance with or without explicit chemistry filters included in the reward signal as a scoring function. MolOpt is the normal benchmark and MolOpt-CF is the benchmark with chemistry filters and a diversity filter applied to each objective.
3.3 Ablation study of the REINVENT algorithm
One of the most popular RL algorithms used for the fine-tuning of CLMs is REINVENT 4. From a theoretical RL perspective, this algorithm can be viewed a combination of the REINFORCE algorithm, experience replay, a reward shaping and a sequence likelihood penalty term for the loss function. However, the impact of each specific mechanism is unclear. Therefore, using the modular components of ACEGEN, we conducted an ablation study to better understand the impact of reward shaping, experience replay, and sequence likelihood penalty, which are key components of the REINVENT algorithm. We used default REINVENT hyperparameters and measured performance on the MolOpt benchmark for comparison. Regarding performance, Table 4 indicates that the primary improvement comes from adding experience replay to REINFORCE. While reward shaping does improve over REINFORCE, combining it with experience replay does not further improve performance. However, adding reward shaping that links the agent policy to the prior policy does improve the chemistry quality, measured by the proportion of valid and unique molecules passing the chemistry filters, albeit with a compromise in performance. These results highlight the importance of regularization in practical applications, as discussed in the previous section. In the original REINVENT implementation, the authors also include a loss term to penalize highly likely sequences. Table 4 shows that this primarily increases exploration but negatively affects all other metrics, including reducing the proportion of molecules passing chemistry filters. Furthermore, if used without reward shaping, it results in a complete collapse in learning.
REINFORCE REINFORCE REINFORCE REINFORCE REINFORCE REINFORCE11footnotemark: 1
- + replay - + replay - + replay
- - + shaping + shaping + shaping + shaping
Metric - - - - + penalty + penalty
Valid 21.74 21.77 21.93 21.90 21.76 21.78
Top-10 Avg 14.74 15.85 14.80 15.69 14.79 15.69
Top-10 AUC 13.21 13.67 13.21 13.57 13.11 13.60
Unique 22.35 22.28 22.46 22.29 22.80 22.63
B&T-CF 14.15 14.34 15.31 15.08 14.83 14.74
B&T-CF Top-10 Avg (Div) 13.81 14.95 13.88 14.80 13.89 14.83
B&T-CF Top-10 AUC (Div) 12.51 12.91 12.48 12.83 12.38 12.85
B&T-CF Diversity (SEDiv@1k) 17.61 17.39 17.93 17.52 18.76 18.19
Table 4: Ablation study of REINFORCE algorithm on the MolOpt benchmark with and without separate components including experience replay, reward shaping, and a high sequence likelihood penalty. Each algorithm was run 5 times with different seeds, and results were averaged and then summed over the 23 tasks. Note that all components together constitute the ACEGEN REINVENT implementation with default hyperparameters.
3.4 Case Study: De-novo generation in the 5-HT2A
To reflect more realistic situations that may arise in practical drug discovery scenarios, we apply the different RL algorithms implemented on a more challenging objective previously proposed by 39. This objective is designing 5-HT2A receptor ligands selective over the highly related D2 receptor utilizing only structural information. This is an important and relevant challenge as clinically, this off-target profile leads to undesirable extrapyramidal side-effects in the treatment of psychosis 48. To achieve this, this objective aims to minimize the docking score against 5-HT2A (PDB: 6A93), yet maximize the docking score against D2 (PDB: 6CM4), within a budget of 32,000 molecules. Both crystal structures are co-crystallized with Risperidone, reflecting the high degree of binding pocket similarity. In contrast to the proposed objective configuration, we use rDock 49 due to more permissive licensing.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Valid 0.95 0.96 0.93 0.94 0.98 0.94 0.95
Top-10 Avg 1.05 1.09 1.10 1.10 1.04 1.06 1.07
Top-10 AUC 1.02 1.05 1.05 1.05 1.02 1.01 1.02
Unique 1.00 1.00 0.89 1.00 1.00 0.99 0.99
B&T-CF 0.71 0.71 0.45 0.63 0.27 0.59 0.48
B&T-CF Top-10 Avg (Div) 1.05 1.06 1.08 1.10 1.02 1.02 1.03
B&T-CF Top-10 AUC (Div) 1.01 1.03 1.04 1.05 1.00 0.99 0.99
B&T-CF Diversity (SEDiv@1k) 0.89 0.91 0.64 0.89 0.77 0.60 0.53
Table 5: Performance summary of algorithms on 5-HT2A case study. Including after applying a basic chemistry filter (B-CF) as described previously. B-CF is the fraction of molecules that pass the chemistry filter. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 32. Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
[…]
Figure 1: Selected examples from the top 10 molecules on the 5HT2A selective task and their docked pose in 5-HT2A (PDB: 6A93). The co-crystallised ligand Risperidone is included as the reference.
3.5 Case study: Scaffold constrained generation
ACEGEN can conduct constrained generation by leveraging PromptSMILES 43 which proposes iterative CLM prompting on SMILES in combination with RL to achieve fine-tune a pre-trained CLMs to the task of constrained generation. Here we demonstrate this application with two experiments: 1) scaffold decoration via known synthetic reactions as proposed with LibINVENT 18, and 2) scaffold decoration followed by scaffold-constrained docking to explore growth vectors inside a binding pocket.
In the first experiment, we compared the performance of RL algorithms on two tasks proposed with LibINVENT 18, both of which conduct scaffold decoration of a piperazine core with an amine linker moiety commonly present in D2 receptor ligands. As in the MolOpt benchmark, we apply a budget of 10,000 molecules. The first task is to optimize the predicted probability of D2 activity as evaluated by a QSAR model. The second task adds a selective reaction filter, rewarding piperazine growth via a Buchwald reaction and amine linker growth via an amide coupling reaction. Table 6 shows successful optimization of both objectives, with AHC performing best on the first task, and REINVENT (either with default or MolOpt parameters) performing best on the second task. However, all algorithms except A2C are able to solve the task with a yield above 90%, and fully satisfy the selective reaction objectives in 80 to 90% of molecules.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Task Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
D2 Yield 0.977 0.989 0.987 0.991 0.354 0.942 0.952
Average score 0.679 0.720 0.743 0.794 0.723 0.749 0.760
D2 with Yield 0.972 0.992 0.988 0.990 0.555 0.950 0.947
reaction Average score 0.579 0.668 0.796 0.723 0.537 0.705 0.726
filters Ratio of satisfied reaction filters 0.896 0.753 0.921 0.809 0.352 0.830 0.858
Table 6: Algorithm comparison in combination with PromptSMILES for constrained molecule generation on LibINVENT DRD2 tasks (without and with selective reaction filters). Each algorithm ran 5 times with different seeds, and results were averaged.
In the second experiment, we used PromptSMILES in combination with AHC (based on the superior performance in the 5-HT2A case study) to conduct scaffold decoration in the catalytic site of BACE1, a key therapeutic target for the treatment of Alzheimer’s disease. Here we used the reference structure 4B05 co-crystallized with AZD3839 50 which is a typical BACE1 inhibitor consisting of an amidine core interacting with two catalytic aspartates, and occupation of the P2’, P1 and P3 sub-pockets 51. The bicyclic, amidine-containing core was used as a prompt for PromptSMILES, allowing two growth vectors, one into the P2’ sub-pocket, and one into the P1 and P3 sub-pockets. Molecules were rewarded by minimizing the docking score using substructure constrained docking, as well as maintaining the number of heavy atoms and rotatable bonds in a sensible range (see Appendix F). Figure 2 shows the successful maximization of the reward over the generation of 10,000 molecules (Figure 2a,b), and corresponding minimization of the docking score (Figure 2c). The top 10 molecules identified contain improved docking scores in the range -23.9 to -15.8 compared to the reference molecule re-docked with a docking score of -8.44. Docked poses (shown in Figure 15) show the recovery of aromatic rings in the P1 and P2 sub-pockets. This case study highlights the successful constrained optimization of a complex objective.
Figure 2: Optimization of the multi-objective reward. The average reward and optimization of the underlying docking score are shown. The top 10 de novo molecules are shown by multi-objective reward, with the constrained substructure highlighted in red and the docking score labeled below. For reference, the co-crystal ligand is re-docked with a docking score of -8.44. The docked poses are shown in Figure 15.
4 Conclusion
In this study, we introduce ACEGEN, a novel toolkit that combines the best methodologies of reinforcement learning (RL) within the field of drug discovery. Leveraging RL building blocks from TorchRL, a general highly-tested decision-making library, ACEGEN provides modular, efficient and versatile solutions for drug discovery that we demonstrate by implementing a suite of language-based solutions. We showcase ACEGEN’s capabilities across diverse areas, including method benchmarking, algorithmic exploration, and real-world drug discovery applications. Our experiments contribute to a better understanding of popular algorithms like REINVENT, offer practical insights into selecting the most suitable algorithm for specific contexts, and underscore the importance of dependable and comprehensive toolkits. Overall, ACEGEN addresses various needs in drug discovery and effectively navigates the complex challenges associated with the field.
Appendix
Appendix A A. Reinforcement Learning Environment
1from acegen.rl_env import SMILESEnv
2from acegen.vocabulary import SMILESVocabulary
3from torchrl.collectors import RandomPolicy
4
5# Create a vocabulary from a list of characters
6chars = ["START", "END", "(", ")", "1", "=", "C", "N", "O"]
7chars_dict = {char: index for index, char in enumerate(chars)}
8vocab = SMILESVocabulary.create_from_dict(chars_dict, start_token="START", end_token="END")
9
10# Create an environment from the vocabulary
11env = SMILESEnv(
12 start_token=vocab.start_token_index,
13 end_token=vocab.end_token_index,
14 length_vocabulary=len(vocab),
15 batch_size=4, # Number of trajectories to collect in parallel
16)
17
18# Create a random policy
19policy = RandomPolicy(env.full_action_spec)
20
21# Take an environment step
22step_tensordict = env.step(policy(env.reset()))
23print(step_tensordict)
[…]
,fontsize=,,
# Output:TensorDict(    fields={        action: Tensor(shape=torch.Size([4]), device=cpu, dtype=torch.int32,        done: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.bool),        next: TensorDict(            fields={                done: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.bool),                observation: Tensor(shape=torch.Size([4]), device=cpu, dtype=torch.int64),                reward: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.float32),                terminated: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.bool),                truncated: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.bool)},            batch_size=torch.Size([4]),            device=cpu,            is_shared=False),        observation: Tensor(shape=torch.Size([4]), device=cpu, dtype=torch.int32),        terminated: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.bool),        truncated: Tensor(shape=torch.Size([4, 1]), device=cpu, dtype=torch.bool)},    batch_size=torch.Size([4]),    device=cpu,    is_shared=False)
Figure 3: Code example that shows how to create a language-based reinforcement learning environment components for SMILES generation and how to take sampling steps on it with a random policy.
Appendix B B. Define Scoring Functions
1from molscore.manager import MolScore
2
3task = MolScore(
4 model_name="example",
5 task_config=path/to/molscore/config/MolOpt/Albyterol_similarity.json,
6 budget=10,
7 output_dir="/tmp",
8))
9
10env = SMILESEnv(
11 start_token=vocab.start_token_index,
12 end_token=vocab.end_token_index,
13 length_vocabulary=len(vocab),
14 batch_size=1,
15)
16
17data = env.rollout(max_steps=100)
18smiles_str = [vocab.decode(smi.numpy()) for smi in data["action"]]
19reward = task(smiles_str)
Figure 4: An example scoring function created using MolScore.
1from rdkit.Chem import AllChem, QED
2
3def evaluate_mol(smiles: str):
4 mol = AllChem.MolFromSmiles(smiles)
5 if mol:
6 return QED(mol)
7 else:
8 return 0.0
9
10def evaluate_mols(smiles: list):
11 return [evaluate_mol(smi) for smi in smiles]
12
13env = SMILESEnv(
14 start_token=vocab.start_token_index,
15 end_token=vocab.end_token_index,
16 length_vocabulary=len(vocab),
17 batch_size=1,
18)
19
20data = env.rollout(max_steps=100)
21smiles_str = [vocab.decode(smi.numpy()) for smi in data["action"]]
22reward = evaluate_mols(smiles_str)
Figure 5: An example of custom scoring function that returns a reward based on an input molecule. If the molecule is invalid, no reward is returned (i.e., 0.0).
Appendix C C. Comparison against an existing implementation
We conducted a comparative analysis between our Reinvent implementation and MolOpt’s 7 implementation, which is accessible on their Github repository. To ensure a fair comparison, we replaced only the prior weights and utilized MolScore for computing scoring functions. Apart from these adjustments, we maintained the use of the original code. Both implementations use the same network architecture and hyperparameters.
Table 7 and Table 8 show that our code is faster and obtains higher values for all metrics with compared to that implementation. The observed performance gaps may stem from subtle idiosyncrasies within the codebase. In reinforcement learning, even minor variations can yield notable discrepancies in outcomes. This underscores the importance of employing standardized, rigorously tested components and foundational building blocks, which offer heightened reliability. Furthermore, we hypothesize the enhanced efficiency can be attributed to leveraging components from TorchRL, a part of the META ecosystem, which inherently provides better optimization and reliability compared to bespoke solutions.
Table 7: Comparison of REINVENT MolOpt Implementation and REINVENT ACEGEN Implementation code on a machine with 32 CPUs and a NVIDIA GeForce RTX 4090 GPU for different scoring functions. In each run the algorithm generates and trains on 10K molecules.
Task REINVENT MolOpt Implementation REINVENT ACEGEN Implementation
Albuterol similarity 2:06 mins 1:01 mins
Amlodipine MPO 5:05 mins 3:25 mins
C7H8N2O2 1:44 mins 0:42 mins
C9H10N2O2PF2Cl 1:44 mins 0:37 mins
Table 8: Comparison of REINVENT MolOpt Implementation and ACEGEN AceGen Implementation on the MolOpt benchmark scoring functions. The results are presented as the sum of all scoring function in the benchmark for different metrics. Each algorithm ran 5 times with different seeds, and results were averaged.
Metric REINVENT MolOpt Implementation REINVENT ACEGEN Implementation
Avg top 10 15.833 15.959
Avg top 100 14.640 14.896
AUC top 10 13.597 13.740
AUC top 100 11.873 11.985
Appendix D D. MolOpt Benchmark
The MolOpt benchmark was run using the MolScore implementation for convenience as it automatically calculates the same metrics as in the original implementation. In addition, MolScore computes the metrics after the removal of undesirable chemistry via the use of a basic chemistry filter (B-CF). This filter ensures molecules have a logP less than or equal to 4.5, rotatable bond count less than or equal to 7, molecular weight in the range 150 to 650 Da, only contain atoms belonging to the following set $A\in\{C,S,O,N,H,F,Cl,Br\}$, and do not violate the substructure alerts as described in 47. If provided with a set of reference molecules, MolScore computes the metrics after the remove of undesirable chemistry via the use of a target chemistry filter (T-CF). This filter ensures molecules have a molecular weight and logP range within $\mu\pm 4\sigma$ of the reference distribution, as well as, removing any molecule that comprises $>10\%$ novel atomic environment bits with respect to the reference molecules, as measured by ECFP4 bits (an example of this is shown in Figure 8). Lastly, MolScore also computes the metrics after filtering molecules using both basic and target chemistry filters (B&T-CF).
Table 9: Performance summary of algorithms for further metrics measured in this benchmark. Each algorithm ran 5 times with different seeds, and results were averaged and then summed over tasks.
AceGen AceGen AceGen AceGen AceGen AceGen AceGen
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Valid 21.77 21.78 21.74 21.45 18.93 21.78 21.52
Top-1 Avg 16.60 16.41 17.73 16.77 16.35 16.14 17.47
Top-10 Avg 15.85 15.69 17.43 16.09 15.67 15.63 17.10
Top-100 Avg 14.59 14.41 16.90 14.95 14.66 14.81 16.50
Top-1 AUC 14.85 14.79 16.28 15.06 15.21 15.39 16.41
Top-10 AUC 13.67 13.60 15.65 13.91 14.27 14.65 15.80
Top-100 AUC 11.96 11.84 14.58 12.17 12.83 13.41 14.67
Top-1 Avg (Div) 16.60 16.41 17.73 16.77 16.35 16.14 17.47
Top-10 Avg (Div) 15.63 15.46 17.13 15.90 15.42 15.35 16.70
Top-100 Avg (Div) 13.89 13.76 14.62 14.21 13.68 12.63 13.54
Top-1 AUC (Div) 14.85 14.79 16.28 15.06 15.21 15.39 16.41
Top-10 AUC (Div) 13.53 13.47 15.42 13.76 14.09 14.37 15.39
Top-100 AUC (Div) 11.66 11.56 13.07 11.80 12.26 11.67 12.36
Unique 22.28 22.63 13.68 22.68 18.38 9.47 10.21
B&T-CF 14.34 14.74 7.00 13.77 7.82 5.81 5.71
B&T-CF Top-1 Avg 15.97 15.86 16.85 16.13 15.61 15.52 16.69
B&T-CF Top-10 Avg 15.21 15.07 16.44 15.44 14.83 14.99 16.32
B&T-CF Top-100 Avg 13.75 13.68 15.61 14.18 13.54 13.98 15.46
B&T-CF Top-1 AUC 14.26 14.20 15.60 14.47 14.54 14.81 15.74
B&T-CF Top-10 AUC 13.06 12.98 14.91 13.27 13.52 14.02 15.11
B&T-CF Top-100 AUC 11.24 11.15 13.61 11.44 11.88 12.64 13.76
B&T-CF Top-1 Avg (Div) 15.97 15.86 16.85 16.13 15.61 15.52 16.69
B&T-CF Top-10 Avg (Div) 14.95 14.83 16.06 15.25 14.54 14.60 15.78
B&T-CF Top-100 Avg (Div) 12.96 12.95 13.12 13.31 12.44 11.41 12.14
B&T-CF Top-1 AUC (Div) 14.26 14.20 15.60 14.47 14.54 14.81 15.74
B&T-CF Top-10 AUC (Div) 12.91 12.85 14.61 13.11 13.32 13.67 14.60
B&T-CF Top-100 AUC (Div) 10.91 10.84 11.84 11.04 11.21 10.59 11.19
B&T-CF Diversity (SEDiv@1k) 17.39 18.19 10.10 17.55 14.24 8.47 7.96
Figure 6: Visualization of the chemical space of the top 100 molecules across all tasks and repetitions by RL algorithm.
Figure 7: Visualization of the consistency of chemical space of the top 100 molecules across all tasks by repetition and RL algorithm.
Figure 8: Demonstration of molecules measured as idiosyncratic by their percentage of novel ECFP4 bits relative to a reference dataset, with novel bits highlighted in red. Molecules are taken from those generated on the JNK3 task and are labelled by RL algorithm, number of novel bits, and percentage of novel bits relative to all bits present in the molecule.
Appendix E E. 5-HT2A case study
[…]
To further probe the differences in the chemistry proposed by the different RL algorithms, the metrics were calculated on all generated de novo molecules within the budget, shown in Tables 12, 13, and 14. Notably, ACEGEN-REINVENT-MolOpt achieved the best average 100 score but performed poorly with respect to chemistry-related metrics such as the ratio of molecules passing quality filters, chemical diversity and idiosyncratic atomic environments as measured by outlier bits with respect to the pre-training dataset. Other algorithms such as, ACEGEN-REINVENT with default parameters, AHC, PPO, and PPOD perform better in these chemistry-related metrics, suggesting an overall higher quality of chemistry proposed. This could reflect stronger regularisation towards the prior distribution, confirmed by higher similarity to the initial pre-training dataset (Table 13). In comparison to the extracted set of known 5-HT2A ligands (Table 14), ACEGEN-AHC manages to generate at least one analog to 78% of known ligands, followed by ACEGEN-REINVENT at 63% and then PPOD at 55%. A visualization of the chemical space explored is shown in 10, showing that ACEGEN-REINFORCE and ACEGEN-REINVENT are most explorative, followed by ACEGEN-AHC, ACEGEN-A2C, ACEGEN-PPO, REINVENT-MolOpt and finally ACEGEN-PPOD least exploratorive. This results in complementary chemical space occupation of the top 100 solutions by either algorithm (Figure 11), all of which show areas of overlap with the known 5-HT2A ligands. Interestingly, a number of the example molecules (Figure 12) contain caged sub-structures or macrocycles typically observed more frequently in natural products than typical small-molecule drugs. This could be a result of the selectivity required in the objective, where more 3D spherical structures are typically more specific binders, but are more challenging synthetically. This can also be seen by a slight shift towards spherical geometries by the ACEGEN-A2C and ACEGEN-PPOD compounds compared to the known selective ligands, shown in Figure 13.
A comparison of the protein interactions made by the top 10 compounds highlight the targeting of potentially selectivity inducing interactions. Figure 14 shows that many of the top 10 compounds proposed by ACEGEN-REINFORCE and ACEGEN-AHC form interactions with S131, T134, W151, and I152 in the second extended binding pocket which differ to those found in D2, and therefore, could potentially be a site to obtain selectivity. For example, occupied by the caged substructure shown in the selected examples in Figure 1 proposed by ACEGEN-AHC, occupying a sub-pocket that is not sterically possible in the D2 crystal structure used here. However, other D2 structures do show occupation of this after re-arrangement of extracellular loop 1 57. This reinforces the difficulty in identifying robust, validated scoring functions or oracles in drug discovery. For example, the use of Gypsum-DL 58 to prepare ligands for docking in this protocol leads to the generation of some implausible tautomers and unlikely protonation states and geometries (for example, cis-amides), meanwhile the docking parameters need further refinement to improve the quality of binding poses and cationic interactions with D1553x32 which are sometimes in an unlikely orientation. These can be swapped for commercial products like LigPrep in MolScore provided the licence is obtained. However, this challenging objective enables the differentiation of RL performance and behaviour in the optimization of a specified objective and sets a realistic and practical benchmark for drug discovery.
Table 10: Description of the scoring functions to define the 5-HT2A selective objective.
Scoring function Comments Target Weight
5-HT2A docking score (inter-molecular only) - Positional cationic restraint for D3x32 - Best score of any protonated/tautomerised/isomeric variant - Normalized based on maximum/minimum values observed at timestep Minimize 1
D2 docking score (inter-molecular only) - Best score of any protonated/tautomerised/isomeric variant - Normalized based on maximum/minimum values observed at timestep Maximize 0.5
Net positive charge 1 1
Maximum number of consecutive rotatable bonds $\leq 3$ 1
Figure 9: Optimisation of a challenging 5-HT2A structure-based selectivity objective by different RL algorithms. The average score does not improve much over the course of training, but the average score of the top 100 compounds improves beyond known selective 5-HT2A ligands. Note that the values here are re-normalized based on the scores achieved for a subset of 5-HT2A ligands that display at least 2-fold selectivity over D2 as extracted from ChEMBL31 32. Therefore, a score of 0.91 indicates the best score observed in the known ligand subset, and a score greater than 1.0 indicates that a single molecule achieves both a better (more negative) 5-HT2A docking score and (more positive) D2 docking score than seen anywhere in the known subset.
Table 11: Performance summary of algorithms for average top-n metrics on 5-HT2A case study. Including after applying a basic chemistry filter (B-CF) as described previously. B-CF is the fraction of molecules that pass the chemistry filter.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Valid 0.95 0.96 0.93 0.94 0.98 0.94 0.95
Top-1 Avg 1.14 1.15 1.20 1.15 1.11 1.09 1.10
Top-10 Avg 1.05 1.09 1.10 1.10 1.04 1.06 1.07
Top-100 Avg 0.97 0.98 1.02 1.02 0.97 0.97 0.98
Top-1 AUC 1.10 1.12 1.12 1.12 1.09 1.06 1.08
Top-10 AUC 1.02 1.05 1.05 1.05 1.02 1.01 1.02
Top-100 AUC 0.93 0.95 0.97 0.96 0.94 0.94 0.94
Top-1 Avg (Div) 1.14 1.15 1.20 1.15 1.11 1.09 1.10
Top-10 Avg (Div) 1.05 1.08 1.10 1.10 1.04 1.06 1.07
Top-100 Avg (Div) 0.97 0.98 1.01 1.02 0.97 0.97 0.98
Top-1 AUC (Div) 1.10 1.12 1.12 1.12 1.09 1.06 1.08
Top-10 AUC (Div) 1.02 1.05 1.05 1.05 1.02 1.01 1.02
Top-100 AUC (Div) 0.93 0.95 0.96 0.96 0.94 0.93 0.94
Unique 1.00 1.00 0.89 1.00 1.00 0.99 0.99
B&T-CF 0.71 0.71 0.45 0.63 0.27 0.59 0.48
B&T-CF Top-1 Avg 1.14 1.13 1.12 1.15 1.11 1.08 1.10
B&T-CF Top-10 Avg 1.05 1.07 1.08 1.10 1.02 1.02 1.03
B&T-CF Top-100 Avg 0.96 0.97 1.00 1.00 0.93 0.95 0.96
B&T-CF Top-1 AUC 1.10 1.11 1.08 1.12 1.09 1.03 1.08
B&T-CF Top-10 AUC 1.01 1.03 1.04 1.05 1.01 0.99 0.99
B&T-CF Top-100 AUC 0.92 0.94 0.96 0.95 0.92 0.93 0.93
B&T-CF Top-1 Avg (Div) 1.14 1.13 1.12 1.15 1.11 1.08 1.10
B&T-CF Top-10 Avg (Div) 1.05 1.06 1.08 1.10 1.02 1.02 1.03
B&T-CF Top-100 Avg (Div) 0.95 0.97 0.98 1.00 0.93 0.95 0.95
B&T-CF Top-1 AUC (Div) 1.10 1.11 1.08 1.12 1.09 1.03 1.08
B&T-CF Top-10 AUC (Div) 1.01 1.03 1.04 1.05 1.00 0.99 0.99
B&T-CF Top-100 AUC (Div) 0.92 0.93 0.94 0.95 0.92 0.92 0.92
B&T-CF Diversity (SEDiv@1k) 0.89 0.91 0.64 0.89 0.77 0.60 0.53
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Validity 0.95 0.96 0.93 0.94 0.98 0.94 0.95
Uniqueness 1.00 1.00 0.89 1.00 1.00 0.99 0.99
Scaffold Uniqueness 0.91 0.90 0.76 0.95 0.98 0.80 0.84
Novelty 1.00 0.99 1.00 1.00 1.00 1.00 1.00
IntDiv 0.87 0.87 0.85 0.86 0.86 0.85 0.83
ScaffDiv 0.85 0.85 0.82 0.85 0.85 0.82 0.81
SEDiv@1k 0.89 0.93 0.64 0.92 0.58 0.70 0.61
SPDiv@1k 0.89 0.89 0.87 0.89 0.87 0.87 0.86
FG 0.02 0.02 0.01 0.03 0.01 0.04 0.05
RS 0.04 0.04 0.03 0.05 0.04 0.02 0.02
Quality Filters 0.85 0.84 0.70 0.84 0.56 0.84 0.78
Predicted Purchasability 0.01 0.02 0.01 0.01 0.01 0.02 0.01
Table 12: Summary of metrics measuring intrinsic properties of the chemistry generated by different algorithms. Most metrics are different measures of chemical diversity, therefore, higher is generally better but can sometimes be a result of idiosyncratic chemistry.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Analogue Similarity 0.28 0.27 0.09 0.21 0.05 0.21 0.22
Analogue Coverage 0.52 0.54 0.21 0.44 0.23 0.27 0.25
FG 0.96 0.98 0.68 0.91 0.69 0.90 0.66
RS 1.00 1.00 0.96 0.98 0.46 0.97 0.97
SNN 0.37 0.37 0.33 0.36 0.27 0.36 0.36
Frag 0.97 0.98 0.89 0.94 0.42 0.96 0.90
Scaf 0.30 0.34 0.07 0.19 0.07 0.14 0.11
OutlierBits 0.08 0.08 0.12 0.10 0.21 0.11 0.12
Table 13: Summary of metrics measuring extrinsic properties of the chemistry generated by different algorithms with respect to the initial training dataset. Metrics are different measures of chemical similarity, therefore, higher is generally better (with the exception of OutlierBits) depending on how much it is desirable to regularize the model and maintain initial prior biases.
ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN ACEGEN
Metric REINFORCE REINVENT REINVENT-MolOpt AHC A2C PPO PPOD
Analogue Similarity 0.01 0.00 0.00 0.00 0.00 0.00 0.01
Analogue Coverage 0.64 0.63 0.39 0.78 0.13 0.25 0.55
FG 0.66 0.65 0.59 0.65 0.42 0.62 0.59
RS 0.75 0.74 0.72 0.73 0.30 0.71 0.76
SNN 0.23 0.22 0.23 0.22 0.18 0.24 0.24
Frag 0.52 0.51 0.59 0.50 0.10 0.50 0.46
Scaf 0.00 0.00 0.00 0.00 0.00 0.00 0.00
OutlierBits 0.41 0.42 0.40 0.43 0.52 0.38 0.39
Table 14: Summary of metrics measuring extrinsic properties of the chemistry generated by different algorithms with respect to known 5-HT2A ligands. Metrics are different measures of chemical similarity, therefore, higher is generally better (with the exception of OutlierBits) depending on how much it is desirable to rediscover known chemistry.
Figure 10: Chemical space explored by different RL algorithms as visualized by UMAP embedding. Known 5-HT2A ligands are plotted in red.
Figure 11: Chemical space of the top 100 molecules identified by each RL algorithms as visualized by UMAP embedding. Known 5-HT2A ligands are plotted in red.
Figure 12: Five example molecules in the range of the top 100 by score. Each row represents REINFORCE, REINVENT, REINVENT-MolOpt, AHC, A2C, PPO and PPOD respectively.
Figure 13: Normalized principal moment ratio space measuring the 3D geometrical shape of the known 5-HT2A ligands and top 100 from generated by the RL algorithms.
Figure 14: Protein ligand interaction fingerprints of the top 10 compounds proposed by each RL algorithm. Examples are selected for visualization of their docking poses in Figure 1. Residue labels marked with an "*" are different for the corresponding structurally aligned residue in D2.
Appendix F F. caffold constrained molecular generation case study
The same training dataset and pre-trained prior was used as for the MolOpt benchmark. In the first experiment, we used the objectives defined by LibINVENT 18 as re-implemented in MolScore. In the second experiment, we conducted a scaffold constrained docking experiment and took the product of three scores in the range [0-1] to define the final reward (see Table 15). The structure used was 4B05 50. Note that during training, the inter-molecular rDock score was used as it was otherwise observed that RL would hack score optimization by improvement of intra-molecular rDock score only (data not shown). To discourage excessive exploitation, a diversity filter was used to penalize the re-generation of molecules via an ’occurance’ filter i.e., up to 5 duplicates resulted in no penalty, 5-10 duplicates resulted in an increasing penalty by multiplication of value scaling linearly from 1 to 0, and more than 10 duplicates resulted in a reward of 0. Both objectives and their configuration files to reproduce them can be found at https://github.com/MorganCThomas/MolScore/.
Scoring function Comments Target
5-HT2A docking score (inter-molecular only) - Substructure restraint based on bicyclic core - Best score of any protonated/tautomerised/isomeric variant at pH 4 - Normalized based on maximum/minimum values observed at timestep Minimize
Maximum number of consecutive rotatable bonds $<3$
Heavy atom count $<50$
Table 15: Description of the scoring functions to define the BACE1 scaffold constrained objective.
Figure 15: The docked poses of (a) AZD3839 re-docked and (b) the top 10 molecules generated by PromptSMILES with AHC. Note that the top 10 molecules often contain an secondary amine in the P1 sub-pocket to form an interaction with Tyr71 and the backbone carbonyls of Phe108 and Lys107 which interact with waters in the co-crystal structure that were removed during preparation. However, none of the top 10 molecules contain a substructure occupying the P3 sub-pocket, as the before mentioned interaction results in an orientation un-amenable to further growth into the P3 sub-pocket, although molecules outside the top 10 do show occupation of the P3 sub-pocket (data not shown).